OpenStudy (anonymous):

the sum 0f 2 numbers is 24 and the first number is 1/5 of the second number

OpenStudy (anonymous):

20 and 4

OpenStudy (anonymous):

I mean 4 and 20

OpenStudy (anonymous):

but they are addition so it doesnt matter..

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

20*.2=4 hence, 20 and 4

OpenStudy (anonymous):

I just guess and plug into calculator and multiply it by one fifth or .2 and add it to the original to see if it's right and I got it on my 3rd try... not really hard

OpenStudy (anonymous):

just have to think a little I guess :/

OpenStudy (bradely):

(1/5)x+x= 24 (2/5)x=24 x=24*5/2 -60 another number is 12 Source: http://www.mathskey.com/question2answer

OpenStudy (kc_kennylau):

(1/5)x+x is not (2/5)x @bradely

OpenStudy (anonymous):

(1/5)x+x.. they are two different numbers. just that one number should be one fifth of another

OpenStudy (anonymous):

sorry Brad

OpenStudy (midhun.madhu1987):